OpenStudy (anonymous):

Factor completely 3x^3 + 9x^2 + x + 3

OpenStudy (ajprincess):

Can u factorise this? 3x^3 + 9x^2 @davidkoveleski

OpenStudy (anonymous):

would it be 12x^5?

OpenStudy (anonymous):

@ajprincess

OpenStudy (ajprincess):

Nope:( \(3x^3=3*x*x*x=3x^2*x\) \(9x^2=3*3*x*x=3x^2*3\) Nw can u pls tell me what is common to both \(3x^3\) and \(9x^2\)?

OpenStudy (anonymous):

3 and x?

OpenStudy (anonymous):

@ajprincess

OpenStudy (ajprincess):

Well actually it is \(3x^2\). \(3x^3=3x^2*x\) and \(9x^2=3x^2*3\). getting it @davidkoveleski

OpenStudy (anonymous):

So how should I answer @ajprincess

OpenStudy (ajprincess):

When we take the common one out what remains is x and 3. So \(3x^3+9x^2=3x^2(x+3)\)

OpenStudy (ajprincess):

x+1=1(x+3) can u factorise \(3x^2(x+3)+1(x+3)\)?

OpenStudy (ajprincess):

u can assume x+3 as y to make factorising simpler. So tell me vat u get when u factorise \(3x^2y+y\). @davidkoveleski

OpenStudy (anonymous):

3x^2y^2 @ajprincess

OpenStudy (ajprincess):

hw did u get it?

OpenStudy (anonymous):

Just put the y's together?

OpenStudy (anonymous):

@ajprincess

OpenStudy (ajprincess):

it can't be done that way.. if it is to do be done there should be * sign between 3x^2y and y. but here u have + sign.

OpenStudy (ajprincess):